Find the amount that results from the given investment.
$300 invested at 9% compounded daily after a period of 2 years
After 2 years, the investment results in $?
(round to the nearest cent as needed)

Answer by mathmate(429)

Question:
Find the amount that results from the given investment.
$300 invested at 9% compounded daily after a period of 2 years

Solution:
APR=9%
daily interest, i = 0.09/365
Principal, P=$300
number of periods (days), n=2*365=730
Future value (amount accumulated)
F=P(1+i)^n
=300(1+.09/365)^730
=300(1.19719)
=$359.16